OpenStudy (buttermequeasy):

Solve for x: |x − 2| + 10 = 12 x = 0 and x = 4 x = −4 and x = 0 x = −20 and x = 4 No Solution

OpenStudy (radar):

|x-2|=12-10 |x-2|=2 x-2=2 and x-2=-2 x=4 x=0